OpenStudy (anonymous):

What is the y coordinate of the y−intercept of the line that passes through the points (−2, −2) and (2, 4)?

Parth (parthkohli):

(x,y) is the basic form. Here the y-coordinates are 4 and -2.

Parth (parthkohli):

-2, 4 in the order.

OpenStudy (anonymous):

And x coordinates are -2, 2.

OpenStudy (anonymous):

3

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

@bitmap Why you say 3?

Parth (parthkohli):

If you wanna determine the equation y = mx+b for two given points, then first determine the slope, then determine the y-intercept by substituting on of the pair of co-ordinates.

OpenStudy (anonymous):

coordinate grid the slope goes through three

OpenStudy (anonymous):

I don't remember the formula to put it into y = mx + b form. :/

OpenStudy (anonymous):

@bitmap Oh, okay.

OpenStudy (anonymous):

i drew a coordinate grid

Parth (parthkohli):

The slope is: \(\Large \color{MidnightBlue}{\Rightarrow {y_{2} - y_{1} \over x_{2} - x_{1}} }\)

Parth (parthkohli):

@bitmap lol no need to do all that.

OpenStudy (anonymous):

I got 2 when I used that formula.

OpenStudy (anonymous):

it was quicker and easier

OpenStudy (anonymous):

it is a small grid

Parth (parthkohli):

Slope is actually 3/2

OpenStudy (anonymous):

ok i suck at math and i tried

OpenStudy (anonymous):

Okay, how would I find the y-coordinate now? @bitmap That's okay. :) So do I.

Parth (parthkohli):

y = (3/2)x + b Substitute y and x with any of the co-ordinates. -2 = (3/2)-2 + b -2 = -3 + b 1 = b

OpenStudy (anonymous):

@ bitmap I appreciate the help you did try to give.

Parth (parthkohli):

So, the equation is y = (3/2)x + 1

Parth (parthkohli):

@bitmap good work :)

Parth (parthkohli):

y intercept is 1.

OpenStudy (anonymous):

Is it 1?

OpenStudy (anonymous):

either 2 or 3

Parth (parthkohli):

y-coordinate? what?

OpenStudy (anonymous):

@ParthKohli Thanks!

Parth (parthkohli):

Haha yw

OpenStudy (anonymous):

Well my problem wanted the y coordinate of the y−intercept. So it is 1?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

Thank. You. So. Much.

Parth (parthkohli):

x-coordinate is 0 y-coordinate is 1.

Parth (parthkohli):

Do you know what an intercept is?

OpenStudy (anonymous):

i tried

OpenStudy (anonymous):

@ParthKohli Yeah. :) @bitmap I know, thank you for trying. :)

OpenStudy (anonymous):

an intercept is where the slope meets the y axis on the graph

Parth (parthkohli):

intercept is the value of the variable when the other variable is 0.

OpenStudy (anonymous):

@ParthKohli Could you help me with this other one? What is the slope of the line passing through the point (4, −2) and perpendicular to the line y = 4x + 2? Write your answer as a fraction.

Parth (parthkohli):

This site explains well :) http://www.purplemath.com/modules/strtlneq3.htm

OpenStudy (anonymous):

Is it the same scenario? And okay, I'll take a look at it. :)

OpenStudy (anonymous):

@ParthKohli and @bitmap Thanks for your help. :)

OpenStudy (anonymous):

your welcome

Parth (parthkohli):

Heehee

OpenStudy (anonymous):

what?